takes the same time to ride each of the miles Does one of the miles appear to have a hil? Mie 1 314 3.25 323 322 321 Mile 2 3.20 322 320 3:16 319 Mile 3 3:34 3 32 3:28 3:32 3:29 Note: when pasting the data into your technology, each mile row will have separate columns for each minute and Print Done c. Reject Ho There is sumcient evidence to warrant rejection of the daim that the three d

media%2F429%2F42925e86-b6c7-4293-b7ba-34
Answer #1

F = 20.9470

p-value = 0.000122

Option A ie Reject H0. Not all Means are equal

Option D ie 3rd Mile has higher mean.

Null and Alternative Hypothesis:

H0: µMile 1 = µMile 2 = µMile 3

H1: Not all Means are equal

Alpha = 0.05

N = 15

n = 5

Degress of Freedom:

dfBetween = a – 1 = 3-1 =2

dfWithin = N-a = 15-3 = 12

dfTotal = N-1 = 15-1 = 14

Critical Values:

Time (dfBetween, dfWithin) : (2,12) = 3.86

Decision Rule:

If F is greater than 3.86, reject the null hypothesis

Test Statistics:

SSBetwen = ?(?ai)2/n - T2/N = 0.03952

SSWithin = ?(Y)2 - ?(?ai)2/n = 0.01132

SSTotal = SSBetwen ­­+ SSWithin = 0.05084

MS = SS/df

F = MSeffect / MSerror

Hence,

F = 0.0198/0.0009 = 20.9470

Result:

Our F = 20.9470, we reject the null hypothesis

Conclusion:

Not all means are equal.